AnimeGhoul8863:

How can the properties of rational exponents be applied to simplify expressions with radicals or rational exponents?

Eiwoh2:

@Elsa213

Elsa213:

@Eiwoh2

Eiwoh2:

Oh come on. x'D

AnimeGhoul8863:

this isnt a call people names post x'D

Eiwoh2:

@Shadow Sorry, but i'm trying to call people who are good at mathetmatics. x'D

AnimeGhoul8863:

x'D its ok

AnimeGhoul8863:

@dude

Elsa213:

https://brainly.com/question/2905460 e.e

dude:

How dare yew e.e

Elsa213:

Dun tell Ultrilliam, boi e.e

Eiwoh2:

x'D

AnimeGhoul8863:

soooooo Dude can yew helps me

dude:

Elsa have you the answer in that link

dude:

gave*

AnimeGhoul8863:

Oh.......ok......it didnt really help but ill read it again

Elsa213:

Breh I just risked my 2% life on this site to answer your question and it didn't help?

Eiwoh2:

Siimplify it, Elsa. x'D Lead the person to the answer, not direct answer it. XD

AnimeGhoul8863:

Sorry elsa but no it was alittle confusing but ill do my best to figure it out

AnimeGhoul8863:

@TheSmartOne can yew help meh

AnimeGhoul8863:

@TheSmartOne

AnimeGhoul8863:

HELP MEHHHHH!

AnimeGhoul8863:

TSO!!!!!!!!

AnimeGhoul8863:

@TheSmartOne PLEASE HELP MEHHHHHHH!

AnimeGhoul8863:

can you make me a radical expression so i can use it to answer this question above

AnimeGhoul8863:

@TheSmartOne THE HECK MAN!

mikewwe13:

The basic property of rational exponents is: a^(1/n) means n√a , that is, the nth root of a. So, for ex, 49^(1/2) means √49, or 7. Also, 8^(1/3) means āˆ›8, or 2. If the numerator is not 1, just use properties of exponents. So, 32^(2/5) = 32^(1/5 * 2) = [ 32^(1/5) ] ^2 = 2^2 = 4